OpenStudy (anonymous):

-6x^2+x+5=0 determine vertex, axis of symmetry, xintercepts; y intercepts, max or min value, domain, range, and end behavior

OpenStudy (triciaal):

one approach set up a table for values of x find the corresponding y values plot the points, sketch the graph and read the results another approach use calculations x- intercept when y =0 factor the function y intercept when x = 0 from the equation in general form ax^2 + bx + c = 0 the x-coordinate of the vertex = -b/2a the y value is the result when the x value is plugged into the function (same as -c/a + b^2/ 4a^2 ) domain = set of x values that satisfies the function range set of y values included

OpenStudy (triciaal):